OpenStudy (anonymous):

A softball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 5. h = height t = time in seconds Find the maximum height of the ball. a) 13.9 feet b) 8.3 feet c) 12.2 feet d) 7.6 feet

OpenStudy (kropot72):

First step is to differentiate with respect to t. Then put the result equal to zero and solve to find the value of t that gives a maximum. Can you differentiate the equation?

OpenStudy (anonymous):

help pls

OpenStudy (anonymous):

maximum height occurs at the vertex... t=-b/(2a)=-32/(2*-16)=1 h(1)= -16(1)^2+32(1)+5=21??

OpenStudy (kropot72):

None of the answers given with the question is correct.

OpenStudy (anonymous):

agree...^^^